OpenStudy (anonymous):

State the property used in each step. Given : 3 + 4(x + 1) 1) Commutative a. = 3 + (4x + 4) 2 )Distributive b. = 4x + (4 + 3) 3) Associative c. = (4x + 4) + 3 4)Addition d. = 4x + 7

OpenStudy (anonymous):

First one is Distributive..

OpenStudy (anonymous):

Distributive Property says that: \[a(b+c) = ab + ac\]

OpenStudy (anonymous):

a) is distributive because it states that a(b+c)=ab+ac. b) is commutative because the terms of the expressions are rearranged and that's what the commutative property states. c) is associative because we are moving the brackets and this property states that (a+b)+c = a+(b+c). d) is addition because 4 is added to 3.

OpenStudy (anonymous):

Commutative property says that: \[a + b = b + a\]

OpenStudy (anonymous):

Associative Property Says that: a + (b + c) = (a + b) + c